OpenStudy (anonymous):

Suppose a magnetic field of 0.20 T is applied to the area within a circular loop of wire. The strength of the field is then changed to 0.10 T. By what factor is the magnetic flux changed? 2 4 1/2 1/4

OpenStudy (anonymous):

IrishBoy123 Are you there?

OpenStudy (irishboy123):

flux = \(\Phi = B \times A\) where B is the magnetic field strength and A is the Area. so, \(\frac{\Phi_2}{\Phi_1} = \frac{0.1T}{0.2T}.\frac{A}{A} = \frac{0.1}{0.2}\) Clearly you need to study the small print : " By what factor is the magnetic flux changed?" http://hyperphysics.phy-astr.gsu.edu/hbase/magnetic/fluxmg.html

OpenStudy (anonymous):

Ok so 0.5 which would then make the answer C. 1/2

OpenStudy (anonymous):

Thank you

OpenStudy (irishboy123):

mp